OpenStudy (anonymous):

10. Twenty-four percent of the patients who take the drug “VIAGRO” suffer from side effects such as runny nose, diarrhea and upset stomach. If n patients at the hospital take VIAGRO, the probability that at most one patient will suffer side effects will be

OpenStudy (anonymous):

what is n?

OpenStudy (anonymous):

the probability that any patient does not suffer is 0.76 the probability that all n patients to not suffer is \[0.76^n\]and the probability that "at least one patient suffers" is 1 - the probability that no patient suffers is \[1-0.76^n\]

OpenStudy (mertsj):

n outcomes in the sample space. 1 successful outcome. Probability = 1/n

OpenStudy (anonymous):

hmm maybe my answer was wrong. at most one means 0 or 1, so you have to compute the probabiity of none, which is \[.76^n\] and the probability of 1 which is \[n\times 00.24\times 0.76^{n-1}\] and add them

OpenStudy (anonymous):

You need to use the following equation where n is the total and r is the number of patients. The probability will be for r=1 and 0 P=nCr (p^r)(p-1)^(n-r) P= P(1) + P(0) = nC1 (0.24)^1 * (0.76)^(n-1) + nC0 (0.24)^0 * (0.76)^(n-0) = 0.24n*(0.76)^(n-1) + 0.76^n

OpenStudy (anonymous):

Sorry the rule I wrote is off P=nCr (p^r)(1-p)^(n-r) It should be 1-p

OpenStudy (anonymous):

right. what zed said. i was wrong and zed is right. it is \[.076^n +n\times 0.24\times 0.76^{n-1}\]

OpenStudy (anonymous):

typo there should be \[0.76^n +n\times 0.24\times 0.76^{n-1}\]